OpenStudy (hellomynameiskam):

What is the sum? 2/3 + 7/15 A. 14/45 B. 1/2 C. 3/5 D. 1 1/15 @rebeccaxhawaii

OpenStudy (alexandervonhumboldt2):

equal the denominators first

OpenStudy (alexandervonhumboldt2):

the least common denominator is 15

rebeccaxhawaii (rebeccaxhawaii):

okay so 3 goes in 15.. 5 times

rebeccaxhawaii (rebeccaxhawaii):

so 10/15 + 7/15

OpenStudy (hellomynameiskam):

17/15

OpenStudy (alexandervonhumboldt2):

\[\frac{ 2 }{ 3 }+\frac{ 7 }{ 15 }=\frac{ 10 }{ 15 }+\frac{ 7 }{ 15 }=\frac{ 10+7 }{ 15 }\]

OpenStudy (alexandervonhumboldt2):

yes

rebeccaxhawaii (rebeccaxhawaii):

that is correct

OpenStudy (alexandervonhumboldt2):

this will be equal to 1 2/15

OpenStudy (alexandervonhumboldt2):

\[1\frac{ 2 }{ 15 }\]

OpenStudy (hellomynameiskam):

okay. Thanks. 2 more

OpenStudy (alexandervonhumboldt2):

ok